OpenStudy (anonymous):

If f(x)=x^2 and g(x)=x+6, find g(f(0))

OpenStudy (usukidoll):

there's another notation for that problem that I think it's easier to see the process g o f this means you put your f(x) function inside of your g(x) whenever you see an x in your function you put that f(x) inside the g(x) then you evaluate when x = 0

OpenStudy (usukidoll):

so my f(x) = x^2 and g(x)=x+6 you put x^2 from f(x) inside the g(x) function (x^2)+6 is your g o f now evaluating at 0 we have (0^2)+6=0+6=6

OpenStudy (usukidoll):

tip: read these types of problems from right to left g o f means place your f(x) function into the g(x) function f o g place your g(x) function into the f(x) function should you have f o f or g o g it means place the f(x) function within itself... similarly place the g(x) within itself if you have a problem like f o g o h you must do g o h first and then use that result to do f o g o h